OpenStudy (dls):

sinx+cosx=1/2 tanx=?

OpenStudy (dls):

\[0<x<\pi \]

OpenStudy (anonymous):

sinx+cosx=1/2 squaring both sides sin^2 x + 2sinx cosx + cos^2 x = 1/4 1 + sin2x = 1/4 sin2x = 1/4 - 1 sin2x = -3/4

OpenStudy (anonymous):

sin2x = 2tanx/ (1+tan^2x) Just find the Value of tanx..) from here

OpenStudy (dls):

why are both of you doing this to me..:'(

OpenStudy (anonymous):

Lol....So...Wat...u Both Two...Approach....) Choose which is easy for u

OpenStudy (dls):

okay..thanks!

OpenStudy (anonymous):

Welcome...:)